# On quasi-Albanese morphisms for log canonical Calabi-Yau pairs

### v1

@@ 3. Proof of the main theorems
Proof of Theorem 2.
Let $\alpha:X\setminus D_{X}\to G$ be the quasi-Albanese morphism. Let $G\subset\mathbb{P}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1})$ be the compactification described in remark 9, where $L_{i}$’s are elements in $\text{Pic}^{0}(A)$ and let $p:\mathbb{P}_{A}\to A$ be the natural projection. Then $H:=\mathbb{P}_{A}\setminus G$ is a snc divisor and $K_{\mathbb{P}_{A}}+H\sim 0$.
Let $\pi:Z\to X$ be a log resolution of $(X,D_{X})$ resolving the indeterminacies of the rational map $X\dashrightarrow\mathbb{P}_{A}$ and let $D_{Z}:=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. By construction, we have
(3) $f^{-1}H\subset D_{Z}.$
Write $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})+E_{Z}$. Since $(X,D_{X})$ is lc, $E_{Z}$ is effective and exceptional over $X$. Then by Theorem 15, $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$, the canonical bundle formula for $(W,D_{W})\to\mathbb{P}_{A}$ writes $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H)$, the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, and $g^{-1}H=D_{W}^{v}$.
[…]
Since the birational map $Z\dashrightarrow W$ contracts precisely the divisor $E_{Z}\subset\mathrm{Ex}(\pi)$, we have $\mu:W\dashrightarrow X$ is a birational contraction and $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$. Let $X\xleftarrow{u}V\xrightarrow{v}W$ be a common log resolution of $(X,D_{X})$ and $(W,D_{W})$. Then $u^{*}(K_{X}+D_{X})=v^{*}(K_{W}+D_{W})$ by the Negativity Lemma, i.e., $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is crepant. By Lemma 6, we have $q(W,D_{W})=q(X,D_{X})$ and $q(W)=q(X)$. By the universal property of quasi-Albanese morphisms and Proposition 12, the composite morphism $W\setminus D_{W}\to W\setminus g^{-1}H\to G$ is the quasi-Albanese morphism of $W\setminus D_{W}$. By the same reasoning, $p\circ g:W\to A$ is the Albanese morphism of $W$. ∎
Theorem 15.
Let $(Z,D_{Z})$ be a projective log smooth pair with $D_{Z}$ reduced and $\kappa(K_{Z}+D_{Z})=0$. Let $A$ be the Albanese variety of $Z$. Let $G$ be a quasi-abelian variety whose abelian part is $A$. Let
$\mathbb{P}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d})$
be a compactification of $G$ where $L_{i}$’s are elements in $\text{Pic}^{0}(A)$. Let $(z_{1},\cdots,z_{d})$ be the inhomogeneous fiber coordinates of $\mathbb{P}_{A}$. For each $i$, let $p_{i}:\mathbb{P}_{A}\to\mathbb{P}_{A}(O_{A}\oplus L_{i})$ be the projection, and let
$H_{i}=(z_{i}=0)+(z_{i}=\infty)\subset\mathbb{P}_{A}(O_{A}\oplus L_{d})$
be the effective and ample over $A$ divisor. Let $H=\sum_{i}p_{i}^{*}H_{i}$.
@@
  • a fibration $f:Z\to\mathbb{P}_{A}$ with $f^{-1}H\subset D_{Z}$,
  • a projective $\mathbb{Q}$-factorial dlt pair $(X,D_{X})$ with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, and
  • a birational morphism $\pi:Z\to X$ such that $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$ and
@@
    Any exceptional over $\mathbb{P}_{A}$ divisor is contained in $E_{Z}+f^{-1}H$.
@@
    the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, and the vertical part $D_{W}^{v}$ of $D_{W}$ coincides with $g^{-1}H$.
@@
Note that by assumption, $G=\mathbb{P}_{A}\setminus H$, $(\mathbb{P}_{A},H)$ is log smooth, and $K_{\mathbb{P}_{A}}+H\sim 0$. Let $F$ be a general fiber of $f$. By Fujino’s subadditivity Theorem 13(3), we have
$0=\nu(K_{Z}+D_{Z})\geq\nu(K_{F}+D_{Z}|_{F})+\kappa(K_{\mathbb{P}_{A}}+H),$
so $\nu(K_{F}+D_{Z}|_{F})=0$. By Gongyo’s abundance theorem [Gon11], $(F,D_{Z}|_{F})$ has a good minimal model and $\kappa(K_{F}+D_{Z}|_{F})=0$.
(1) Let $Q\subset Z$ be an exceptional over $\mathbb{P}_{A}$ prime divisor which is not contained in $f^{-1}H$. Let $q:\mathbb{P}^{\prime}_{A}\to\mathbb{P}_{A}$ be a composition of the blowup along $f(Q)$ with a log resolution. We may assume that $q$ is an isomorphism over $\mathbb{P}_{A}\setminus f(Q)$. Let $p:Z^{\prime}\to Z$ be a log resolution of $(Z,D_{Z})$ resolving the indeterminacies of $Z\dashrightarrow\mathbb{P}_{A}^{\prime}$. We have the following commutative diagram.
@@
Write $K_{Z^{\prime}}+D_{Z^{\prime}}=p^{*}(K_{Z}+D_{Z})+E_{1}$, $K_{\mathbb{P}^{\prime}_{A}}+H^{\prime}=q^{*}(K_{\mathbb{P}_{A}}+H)+E_{2}$, where $H^{\prime}=q^{-1}H$, $D_{Z^{\prime}}=p^{-1}_{*}D_{Z}+\operatorname{Ex}(p)$, and $E_{1}$, $E_{2}$ are effective and exceptional divisors. By $f^{\prime-1}H^{\prime}\subset D_{Z^{\prime}}$,
@@
and Theorem 13(1), we have $E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z}$ is pseudoeffective. Since $(\pi\circ p)_{*}(E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z})=-(\pi\circ p)_{*}f^{\prime*}E_{2}$ is pseudoeffective, we have $f^{\prime*}E_{2}$ is exceptional over $X$. Hence $E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z}$ is effective by a Lemma due to Lazarsfeld (cf. [KL09, Corollary 13]).
@@
$f^{-1}(P+H)\subset E_{Z}+f^{-1}H\subset D_{Z}.$
Then by Lemma 14 and Lemma 16, we have
$0=\nu(K_{Z}+D_{Z})=\overline{\nu}(Z\setminus D_{Z})\geq\overline{\nu}(F\setminus D_{Z})+\overline{\kappa}({\mathbb{P}_{A}}\setminus P+H)=\overline{\kappa}(G\setminus P)>0,$
@@
(3) By (2) and the assumption in (3), we known that the vertical part $E_{Z}^{v}$ of $E_{Z}$ is very exceptional over $\mathbb{P}_{A}$. Since $\kappa(E_{Z}|_{F})=\kappa(K_{F}+D_{Z}|_{F})=0$, we have
(4) $(f_{*}O_{Z}(mE_{Z}))^{\wedge}=O_{\mathbb{P}_{A}}$
@@
for all sufficiently divisible $m>0$, where $\wedge$ means taking reflexive hull.
By Gongyo’s abundance theorem and Hacon-Xu [HX13, Theorem 2.12], $(Z,D_{Z})$ has a good minimal model over a nonempty Zariski open subset $U$ of $\mathbb{P}_{A}$. We run a $(Z,D_{Z})$-MMP over $\mathbb{P}_{A}$ with scaling of an ample divisor. Then this MMP will terminate over $U$ by [HX13, Corollary 2.9]. Hence we have a partial MMP $(Z,D_{Z})\dashrightarrow(Z^{\prime},D_{Z^{\prime}})$ with $K_{Z^{\prime}}+D_{Z^{\prime}}$ semiample over $U$, where $D_{Z^{\prime}}$ is the pushforward of $D_{Z}$. Let $F^{\prime}$ be a general fiber of $f^{\prime}:Z^{\prime}\to\mathbb{P}_{A}$ and let $E_{Z^{\prime}}$ be the pushforward of $E_{Z}$. Since $\kappa(K_{F^{\prime}}+D_{Z^{\prime}}|_{F^{\prime}})=0$, we have $E_{Z^{\prime}}|_{F^{\prime}}\sim_{\mathbb{Q}}K_{Z^{\prime}}+D_{Z^{\prime}}|_{F^{\prime}}\sim_{\mathbb{Q}}0$, so $E_{Z^{\prime}}$ is vertical over $\mathbb{P}_{A}$. By formula (4), we have $(f^{\prime}_{*}O_{Z^{\prime}}(mE_{Z^{\prime}}))^{\wedge}=(f_{*}O_{Z}(mE_{Z}))^{\wedge}=O_{\mathbb{P}_{A}}$ for all sufficiently divisible $m>0$. Hence $E_{Z^{\prime}}$ is very exceptional very $\mathbb{P}_{A}$ by [GL13, Lemma 2.15]. Then $(Z^{\prime},D_{Z^{\prime}})$, and so $(Z,D_{Z})$, has a good minimal model over $\mathbb{P}_{A}$ by Birkar [Bir12, Theorem 1.8].
@@
(4) Write the canonical bundle formula for $g:(W,D_{W})\to\mathbb{P}_{A}$ as $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+B_{\mathbb{P}_{A}}+M_{\mathbb{P}_{A}})$. Since $f^{-1}H\subset D_{Z}$, we have $g^{-1}H\subset D_{W}$, so $B_{\mathbb{P}_{A}}\geq H$. Since $K_{W}+D_{W}\sim_{\mathbb{Q}}0$ and $K_{\mathbb{P}_{A}}+H\sim 0$, we have $B_{\mathbb{P}_{A}}=H$ and $M_{\mathbb{P}_{A}}\sim_{\mathbb{Q}}0$, so the canonical bundle formula writes
(5) $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H).$
@@
Let $\Sigma$ be a reduced snc divisor on $G$ and let $\overline{\Sigma}$ be its closure in $\mathbb{P}_{A}$. Then the canonical bundle formula for $(W,D_{W}+g^{*}\overline{\Sigma})\to\mathbb{P}_{A}$ writes $K_{W}+D_{W}+g^{*}\overline{\Sigma}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H+\overline{\Sigma})$. By inversion of adjunction (cf. [Fil20, Proposition 4.16]), $(W,D_{W}+g^{*}\overline{\Sigma})|_{g^{-1}G}$ is lc. Hence the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable by [Kol23, Corollary 4.55]. In particular, the restriction $g|_{g^{-1}G}$ is equidimensional, and so $W$ has no exceptional divisors over $G$.
Let $Q\subset D_{W}^{v}$ be a prime divisor. If $g(Q)\not\subset H$, then $g(Q)$ is a divisor on $\mathbb{P}_{A}$. But by the canonical bundle formula (5), we must have $H\geq g(Q)$, which is a contradiction. Hence $D_{W}^{v}=g^{-1}H$.
(5) By (3), it suffices to show that for any prime divisor $P\subset H$, there is a prime divisor $Q\not\subset E_{Z}$ dominating $P$. Assume the contrary, i.e., assume that there is a prime divisor $P\subset H$, and an effective exceptional over $\mathbb{P}_{A}$ divisor $E$ with $f(E)\subset P$ such that
(6) $f^{-1}P\subset E_{Z}+E.$
Case 1. $\dim\mathbb{P}_{A}-\dim A=1$. By [BFPT24, Theorem 1.5], the Albanese morphism $p_{X}:X\to A^{\prime}$ of $X$ is a fibration. Since $X$ has rational singularities, we have $\dim A^{\prime}=q(X)=q(Z)=\dim A$. Hence, the induced morphism $A\to A^{\prime}$ is an isomorphism. We identify $A^{\prime}$ with $A$. We have the following commutative diagram.
@@
where $p$ is the natural projection.
We claim that
(7) $f^{-1}P\subset E_{Z}.$
Let $Q\subset E$ be a prime divisor. Then $Q\subset f^{-1}H\subset D_{Z}$. Since $\dim\mathbb{P}_{A}-\dim A=1$, $f(Q)$ is vertical over $A$, so $f(Q)$ is not an lc center of $(X,D_{X})$ by [BFPT24, Theorem 1.5]. Hence
$1<a(Q,X,D_{X})=-\operatorname{mult}_{Q}(D_{Z}-E_{Z})=-1+\operatorname{mult}_{Q}E_{Z}.$
Hence $Q\subset E_{Z}$ and the claim follows.
Let $L_{A}$ be an effective and sufficiently ample divisor on $A$. Since $P$ is ample over $A$, we have
$\dim\mathbb{P}_{A}$ $=\kappa(P+p^{*}L_{A})$
$=\kappa(f^{*}P+(p\circ f)^{*}L_{A})$
[…]
$=\kappa(E_{Z}+(p_{X}\circ\pi)^{*}L_{A})$
$=\kappa(p_{X}^{*}L_{A})\text{ ($E_{Z}$ is exceptional over $X$)}$
$=\dim A,$
which is a contradiction, so the inclusion (6) can not be true.
Case 2. $\dim\mathbb{P}_{A}-\dim A>1$. Since $P\subset H=\sum_{i}p_{i}^{*}H_{i}$, we have $P\subset p_{i}^{*}H_{i}$ for some $i$. By re-indexing $i$’s, we may assume that $P\subset p_{1}^{*}H_{1}$. Let
$\mathbb{P}^{\prime}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{2})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}),$
let $p_{i}^{\prime}:\mathbb{P}_{A}^{\prime}\to\mathbb{P}_{A}(O_{A}\oplus L_{i})$, $2\leq i\leq d$, $\alpha:\mathbb{P}_{A}\to\mathbb{P}_{A}^{\prime}$ be the natural projections and let $H^{\prime}=\sum_{i=2}^{d}{p^{\prime}}_{i}^{*}H_{i}$. Then
$(\alpha\circ f)^{-1}H^{\prime}\subset f^{-1}H\subset D_{Z}.$
@@
where $p^{\prime}$ is the natural projection and $p_{V}$ is the Albanese morphism of $V$ (by the same reasoning as in Case 1, we may identify the Albanese varieties of $Z$ and $V$).
@@
(8) $f^{-1}P\subset E_{Z}+(\alpha\circ f)^{-1}H^{\prime}.$
Let $Q\subset E$ be a prime divisor which is not contained in $E_{Z}$. Then $Q\subset f^{-1}H\subset D_{Z}$. Since $Z\dashrightarrow V$ contracts precisely the divisor $E_{Z}$, the strict transform $Q_{V}$ of $Q$ on $V$ is a divisor contained in $D_{V}$. Since $\dim\mathbb{P}_{A}-\dim\mathbb{P}_{A}^{\prime}=1$, and $Q$ is exceptional over $\mathbb{P}_{A}$, we have $Q$ is vertical over $\mathbb{P}_{A}^{\prime}$. Hence $Q_{V}$ is vertical over $\mathbb{P}_{A}^{\prime}$. By (4), we have $g(Q_{V})\subset H^{\prime}$. Hence $Q\subset(\alpha\circ f)^{-1}H^{\prime}$ and the claim follows.
@@
which is a contradiction, so the inclusion (6) can not be true. ∎
Lemma 16.

### v2

@@ 3. Proof of the main theorems
3.1. Proof of theorem 1.2
Proof of theorem 1.2.
This is (3) of the following Theorem 3.1. ∎
Theorem 3.1.
  1. (1)
    Let $(X^{0},D_{X}^{0})$ be a quasi-projective log smooth lc pair with $\overline{\kappa}(X^{0},D_{X}^{0})=0$. Then the quasi-Albanese morphism of $X^{0}$ is a fibration.
  2. (2)
    Let $(X,D_{X})$ be a projective $\mathbb{Q}$-factorial dlt pair with $\kappa(K_{X}+D_{X})=0$. Let $0\leq\Delta_{X}\leq\lfloor D_{X}\rfloor$ be a reduced divisor. Then the quasi-Albanese morphism of $X\setminus\Delta_{X}$ is a fibration.
  3. (3)
    Let $(X,D_{X})$ be a projective lc pair with $\kappa(K_{X}+D_{X})=0$. Then the quasi-Albanese morphism of $X\setminus\operatorname{nklt}(X,D_{X})$ is a fibration.
Proof.
(1) The proof is identical to that of [fujino2024quasi, Theorem 1.2] by using Lemma 2.9 (1) instead of [fujino2024quasi, Theorem 6.1].
(2) Let $\pi:Z\to X$ be a log resolution and let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Since $\pi^{-1}\Delta_{X}\leq\lfloor D_{Z}\rfloor$, we have $\overline{\kappa}(Z\setminus\pi^{-1}\Delta_{X},D_{Z}-\pi^{-1}\Delta_{X})=\kappa(K_{Z}+D_{Z})=\kappa(K_{X}+D_{X})=0$. By (1), the quasi-Albanese morphism of $Z\setminus\pi^{-1}\Delta_{X}$ is a fibration. Hence we have (2).
(3) Let $\pi:Z\to X$ be a log resolution such that $\pi^{-1}\operatorname{nklt}(X,D_{X})$ is a divisor. Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Since $\pi^{-1}\operatorname{nklt}\leq\lfloor D_{Z}\rfloor$, we have $\overline{\kappa}(Z\setminus\pi^{-1}\operatorname{nklt},D_{Z}-\pi^{-1}\operatorname{nklt})=\kappa(K_{Z}+D_{Z})=\kappa(K_{X}+D_{X})=0$. By (1), the quasi-Albanese morphism of $Z\setminus\pi^{-1}\operatorname{nklt}$ is a fibration. Hence we have (3). ∎
3.2. Proof of theorem 1.4
In the following theorem, we do not assume that $G$ is the quasi-Albanese variety of $Z\setminus\lfloor D_{Z}\rfloor$ and we do not assume that $G\neq A$ (in particular, $(\mathbb{P}_{A},H)$ may coincides with $(A,0)$).
Theorem 3.2.
Let $(Z,D_{Z})$ be a projective log smooth lc pair with $\kappa(K_{Z}+D_{Z})=0$. Let $G$ be a quasi-abelian variety and let $A$ be its compact part. Let $(\mathbb{P}_{A},H)$ be the canonical compactification of $G$. Note that $(\mathbb{P}_{A},H)$ is log smooth and $K_{\mathbb{P}_{A}}+H\sim 0$.
@@
  • a fibration $f:Z\to\mathbb{P}_{A}$ with $f^{-1}H\subset\lfloor D_{Z}\rfloor$,
  • a projective lc pair $(X,D_{X})$ with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, and
  • a birational morphism $\pi:Z\to X$ such that $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Note that since $(X,D_{X})$ is lc, we have
@@
    Let $Q\subset Z$ be a prime divisor which is exceptional over $\mathbb{P}_{A}$. Then $Q\subset E_{Z}+f^{-1}H$.
@@
    the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, and $D_{W}^{v}=\lfloor D_{W}^{v}\rfloor=g^{-1}H$. The induced birational map $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is a crepant birational contraction with $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$.
@@
Let $F$ be a very general fiber of $f$. By Fujino’s subadditivity Theorem 2.8(3), we have
$0=\nu(K_{Z}+D_{Z})\geq\nu(K_{F}+D_{Z}|_{F})+\kappa(K_{\mathbb{P}_{A}}+H).$
Hence $\nu(K_{F}+D_{Z}|_{F})=0$. By Gongyo’s abundance theorem [gongyo2011minimal], we have $\kappa(K_{F}+D_{Z}|_{F})=0$ and $(F,D_{Z}|_{F})$ has a good minimal model.
(1) Assume that $Q\not\subset f^{-1}H$. Let $q:\mathbb{P}^{\prime}_{A}\to\mathbb{P}_{A}$ be a composition of the blowup along $f(Q)$ with a log resolution. We may assume that $q$ is an isomorphism over $\mathbb{P}_{A}\setminus f(Q)$. Let $p:Z^{\prime}\to Z$ be a log resolution of $(Z,D_{Z})$ resolving the indeterminacies of $Z\dashrightarrow\mathbb{P}_{A}^{\prime}$. We have the following commutative diagram.
@@
Write $K_{Z^{\prime}}+D_{Z^{\prime}}=p^{*}(K_{Z}+D_{Z})+E_{1}$, $K_{\mathbb{P}^{\prime}_{A}}+H^{\prime}=q^{*}(K_{\mathbb{P}_{A}}+H)+E_{2}$, where $H^{\prime}=q^{-1}H$, $D_{Z^{\prime}}=p^{-1}_{*}D_{Z}+\operatorname{Ex}(p)$, and $E_{1}$, $E_{2}$ are effective and exceptional divisors. By $f^{-1}H\subset\lfloor D_{Z}\rfloor$, we have $f^{\prime-1}H^{\prime}\subset\lfloor D_{Z^{\prime}}\rfloor$. By
@@
$\kappa(K_{F}+D_{Z}|_{F})=0$, and Theorem 2.8(2), we have $E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z}$ is pseudo-effective. Since $(\pi\circ p)_{*}(E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z})=-(\pi\circ p)_{*}f^{\prime*}E_{2}$ is pseudo-effective, we have $f^{\prime*}E_{2}$ is exceptional over $X$. Hence $E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z}$ is effective by a Lemma due to Lazarsfeld (cf. [kollar2010quotients, Corollary 13]).
@@
$f^{-1}(P+H)\subset E_{Z}+f^{-1}H\subset\lfloor D_{Z}\rfloor.$
Consider the fibration $Z\setminus\lfloor D_{Z}\rfloor\to\mathbb{P}_{A}\setminus P+H$. By Lemma 2.9(2) and Lemma 3.2, we have
$0=\nu(K_{Z}+D_{Z})=\overline{\nu}(Z\setminus\lfloor D_{Z}\rfloor,D_{Z})\geq\overline{\nu}(F\setminus\lfloor D_{Z}\rfloor,D_{Z})+\overline{\kappa}({\mathbb{P}_{A}}\setminus P+H)=\overline{\kappa}(G\setminus P)>0,$
@@
(3) By (2) and the assumption in (3), we known that the vertical part $E_{Z}^{v}$ of $E_{Z}$ is degenerate over $\mathbb{P}_{A}$. Since $\kappa(E_{Z}|_{F})=\kappa(K_{F}+D_{Z}|_{F})=0$, we have
(3.2.1) $O_{\mathbb{P}_{A}}\subset f_{*}O_{Z}(mE_{Z})\subset(f_{*}O_{Z}(mE_{Z}))^{\wedge}=O_{\mathbb{P}_{A}}$
@@
for all sufficiently divisible $m>0$, where $\wedge$ means taking reflexive hull. In particular,
$\bigoplus_{m\geq 0}f_{*}O_{Z}(m(K_{Z}+D_{Z}))=\bigoplus_{m\geq 0}f_{*}O_{Z}(mE_{Z})$
is a finitely generated $O_{\mathbb{P}_{A}}$-algebra. By Gongyo’s abundance theorem and Hacon-Xu [hacon2013existence, Theorem 2.12], $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$. Denote by $g:W\to\mathbb{P}_{A}$ the induced fibration. Let $m>0$ be a sufficiently divisible integer. Then there is a surjective morphism
$g^{*}g_{*}O_{W}(m(K_{W}+D_{W}))\to O_{W}(m(K_{W}+D_{W})).$
By formula (3.2.1), we have
$g_{*}O_{W}(m(K_{W}+D_{W}))=f_{*}O_{Z}(m(K_{Z}+D_{Z}))=O_{\mathbb{P}_{A}}.$
Hence $K_{W}+D_{W}\sim_{\mathbb{Q}}0$.
@@
(4) Write the canonical bundle formula for $g:(W,D_{W})\to\mathbb{P}_{A}$ as $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+B_{\mathbb{P}_{A}}+M_{\mathbb{P}_{A}})$. Since $f^{-1}H\subset\lfloor D_{Z}\rfloor$, we have $g^{-1}H\subset\lfloor D_{W}\rfloor$. Hence $B_{\mathbb{P}_{A}}\geq H$. Since $K_{W}+D_{W}\sim_{\mathbb{Q}}0$ and $K_{\mathbb{P}_{A}}+H\sim 0$, we have $B_{\mathbb{P}_{A}}=H$ and $M_{\mathbb{P}_{A}}\sim_{\mathbb{Q}}0$. Hence the canonical bundle formula writes
(3.2.2) $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H).$
@@
Let $\Sigma$ be a reduced snc divisor on $G$ and let $\overline{\Sigma}$ be its closure in $\mathbb{P}_{A}$. Then the canonical bundle formula for $(W,D_{W}+g^{*}\overline{\Sigma})\to\mathbb{P}_{A}$ writes $K_{W}+D_{W}+g^{*}\overline{\Sigma}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H+\overline{\Sigma})$. By inversion of adjunction (cf. [Ambro2004Shokurov, Theorem 3.1] or [filipazzi2018generalized, Proposition 4.16]), $(W,D_{W}+g^{*}\overline{\Sigma})|_{g^{-1}G}$ is lc. Hence the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable by [Kollar-Families, Corollary 4.55]. In particular, by [Kollar-Families, Definition-Theorem 4.7], the restriction $g|_{g^{-1}G}$ is flat.
Let $Q\subset D_{W}^{v}$ be a prime divisor. If $g(Q)\not\subset H$, then $P:=g(Q)$ is a divisor on $\mathbb{P}_{A}$. Since $(W,D_{W}+g^{*}P)$ is lc near the generic point of $P$, we have
$1\leq a(Q,W,D_{W}+g^{*}P)=-\operatorname{mult}_{Q}(D_{W}+g^{*}P)=-1-\operatorname{mult}_{Q}D_{W}<-1,$
which is a contradiction. Hence $D_{W}^{v}\subset g^{-1}H\subset\lfloor D_{W}^{v}\rfloor$.
Since the birational map $Z\dashrightarrow W$ contracts precisely the divisor $E_{Z}\subset\mathrm{Ex}(\pi)$, we have $W\dashrightarrow X$ is a birational contraction and $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$. Let $X\xleftarrow{u}V\xrightarrow{v}W$ be a common log resolution of $(X,D_{X})$ and $(W,D_{W})$. Then $u^{*}(K_{X}+D_{X})=v^{*}(K_{W}+D_{W})$ by the Negativity Lemma (cf. [KM98, 3.39]), i.e., $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is crepant.
(5) Case 1. $\dim\mathbb{P}_{A}-\dim A=0$. By Gongyo’s abundance theorem, $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$. Let $Z\xleftarrow{u}V\xrightarrow{v}W$ be a common resolution. By [hacon2007shokurov], the rigidity Lemma, and the universal property for Albanese morphisms, we have the following commutative diagram.
@@
where $p_{Z}$, $p_{W}$, $p_{Z}\circ u$, $p_{W}\circ v$ are Albanese morphisms, $A_{Z}\to A$ is the induced morphism by the universal property of $p_{Z}$. Hence we may identify the Albanese varieties of $W$ and $Z$, and there is a morphism $W\to A$, compatible with the rational map $Z\dashrightarrow W$. Hence $(W,D_{W})$ is actually a good minimal model of $(Z,D_{Z})$ over $A$.
Case 2. $\dim\mathbb{P}_{A}-\dim A>0$. By (3),to prove (5), it suffices to show that for any prime divisor $P\subset H$, there is a prime divisor $Q\not\subset E_{Z}$ dominating $P$. Assume the contrary, i.e., assume that there is a prime divisor $P\subset H$, and an effective exceptional over $\mathbb{P}_{A}$ divisor $E$ with $f(E)\subset P$ such that
(3.2.3) $f^{-1}P\subset E_{Z}+E.$
Let $L_{i}\in\text{Pic}^{0}(A)$, $1\leq i\leq d$, be line bundles such that $G$ is the fiber product over $A$ of the total spaces of $L_{i}$’s with their zero sections removed. Then $\mathbb{P}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1})$. Let $(z_{1},\cdots,z_{d})$ be the inhomogeneous fiber coordinates of $\mathbb{P}_{A}$. For each $i$, let $p_{i}:\mathbb{P}_{A}\to\mathbb{P}_{A}(O_{A}\oplus L_{i}^{-1})$ be the projection, and let
$H_{i}=(z_{i}=0)+(z_{i}=\infty)\subset\mathbb{P}_{A}(O_{A}\oplus L_{i}^{-1})$
be the effective and ample over $A$ divisor. Then $H=\sum_{i=1}^{d}p_{i}^{*}H_{i}$.
Since $P\subset H=\sum_{i=1}^{d}p_{i}^{*}H_{i}$, we have $P\subset p_{i}^{*}H_{i}$ for some $i$. By re-indexing $i$’s, we may assume that $P\subset p_{1}^{*}H_{1}$. Let
$\mathbb{P}^{\prime}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{2}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1}),$
and let $p_{i}^{\prime}:\mathbb{P}_{A}^{\prime}\to\mathbb{P}_{A}(O_{A}\oplus L_{i}^{-1})$, $2\leq i\leq d$, $\alpha:\mathbb{P}_{A}\to\mathbb{P}_{A}^{\prime}$, $p^{\prime}:\mathbb{P}_{A}^{\prime}\to A$ be the natural projections. Let $H^{\prime}=\sum_{i=2}^{d}{p^{\prime}}_{i}^{*}H_{i}$. Then $\mathbb{P}_{A}^{\prime}\setminus H^{\prime}$ is a quasi-abelian variety whose canonical compactification is just $(\mathbb{P}_{A}^{\prime},H^{\prime})$, and $\alpha\circ f:Z\to\mathbb{P}_{A}^{\prime}$ is a fibration with
$(\alpha\circ f)^{-1}H^{\prime}\subset f^{-1}H\subset\lfloor D_{Z}\rfloor.$
@@
(3.2.4) $f^{-1}P\subset E_{Z}+(\alpha\circ f)^{-1}H^{\prime}.$
Let $Q\subset E$ be a prime divisor which is not contained in $E_{Z}$. Since $f(E)\subset P\subset H$, we have $Q\subset f^{-1}H\subset\lfloor D_{Z}\rfloor$. Since the MMP $Z\dashrightarrow V$ contracts precisely the divisor $E_{Z}$, the strict transform $Q_{V}$ of $Q$ on $V$ is a divisor contained in $\lfloor D_{V}\rfloor$. Since $\dim\mathbb{P}_{A}-\dim\mathbb{P}_{A}^{\prime}=1$, and $Q$ is exceptional over $\mathbb{P}_{A}$, we have $Q$ is vertical over $\mathbb{P}_{A}^{\prime}$. Hence $Q_{V}$ is vertical over $\mathbb{P}_{A}^{\prime}$. By (4), we have $Q_{V}\subset g^{-1}H^{\prime}$. Hence $Q\subset(\alpha\circ f)^{-1}H^{\prime}$ and the claim follows.
@@
which is a contradiction. Hence the inclusion (3.2.3) can not be true. ∎
Lemma 3.3 (cf. [Deng2024HyperFund, Proposition 1.3]).